# Control the GNN: Utilizing Neural Controller with Lyapunov Stability for Test-Time Feature Reconstruction

### v1

@@ Introduction
where $Y$ is the equilibrium point and $\epsilon$ is a small positive number. The condition $V_{\phi}(\hat{Y})\leq 0\cup\Delta V_{\phi}(\hat{Y})\geq 0$ restricts the found state $\hat{y}$ from satisfying the conditions in Equation 7. The constraint $\lVert\hat{Y_{i}}-Y_{i}\rVert_{2}\geq\epsilon$ ensures numerical stability. When the state is close to the equilibrium point, numerical calculations may become unstable, especially when using solvers to analyze system dynamics. By introducing this term, we ensure that the numerical solution does not produce large relative errors or numerical overflow when the state is close to zero. We utilize the satisfiability modulo theories (SMT) solver to solve this equation, obtaining states that violate the Lyapunov criterion. When the solver finds states that violate the Lyapunov criterion within the domain, we add these states as counterexamples to the training set for training the neural controller $f_{\theta}$ and the neural Lyapunov function $V_{\phi}$. Our training process is detailed in Algorithm 1.
@@
16:     Prepare close-loop system $C$ and $\hat{Y}$ derived from well-trained GNN
@@
Test-Time Feature Reconstruction
@@
Table 1: Results(%) for the semi-supervised node classification task on the four datasets. Training split represents the method for selecting training nodes. ’Bias’ means that training nodes are selected using the Scalable Biased Sampler (Zhu et al. 2021). Biased training samples simulate the situation with distribution shift. Each result is reported as an average $\pm$ standard deviation across 10 experiments. The footnote of FRGNN indicates the base model of FRGNN. * indicates that the FRGNN only replaces one class of node features. OOM indicates that the method runs out of memory. The best results are highlighted in bold and the second-best results are underlined.
@@
Experimental Setup
Base Model
We chose SGC  (Wu et al. 2019) as our base model. SGC is a simple graph neural network that reduces the multi-layer convolution operations of GCN  (Kipf and Welling 2017) to single matrix multiplication, making it more efficient. Our method requires using an SMT solver to identify states that violate Lyapunov constraints, so we must keep the SMT solving time within an acceptable range. Thus, we select SGC as the base model to accelerate the solving process by reducing the graph neural network’s parameters.
Datasets
We evaluate our method on five popular graph datasets: Cora, Citeseer, Pubmed (Yang, Cohen, and Salakhutdinov 2016), Amazon Photo, and Amazon Computers (Shchur et al. 2019). These datasets are standard benchmarks for semi-supervised node classification tasks. The statistics of these datasets are summarized in Appendix C. For each dataset, we follow the setup from FRGNN  (Ding et al. 2024) and apply the Scalable Biased Sampler (Zhu et al. 2021) to create training, validation, and test sets. This sampling method effectively captures datasets with distribution shifts, allowing us to demonstrate the ability of our method to address such issues.
Table 2: Classification accuracy comparison between our neural controller and traditional controller on three datasets. Each result is reported as an average $\pm$ standard deviation across 10 experiments. MFAC indicates that the controller is MFAC and NC indicates that the controller is our neural controller. The best results are highlighted in bold.
@@
Implementation Details
Our training process for the neural controller and the neural Lyapunov function heavily relies on generating counterexamples, which necessitates ensuring the speed of the SMT solver. Since the speed of SMT solver is proportional to the number of parameters in each graph neural network layer, we downsample the node features to reduce the parameter count. We apply PCA to reduce the feature dimensions to a fixed 20 dimensions, making it the input size for all experiments. We use SGC as the base model, setting its steps of feature propagation to 3 and its output size to the number of classes. Our neural controller $f_{\theta}$ is a two-layer MLP with a hidden size of 16. Our neural Lyapunov function $V_{\phi}$ is also a two-layer MLP with a hidden size of 16 and an output size of 1. We train both $f_{\theta}$ and $V_{\phi}$ using the Adam optimizer with a learning rate of 0.001. We conduct all experiments on an RTX 3090 24G GPU. More details are provided in the Appendix B.
@@
Baseline
We select MH-AUG  (Park et al. 2021), KDGA  (Wu et al. 2022), and FRGNN  (Ding et al. 2024) as the baselines. More detail about these methods is provided in the Appendix D.
@@
Results
Performance Comparison

### v2

@@ III-C Training Methodology
where $Y$ is the equilibrium point and $\epsilon$ is a small positive number. We use the SMT solver to find the counterexamples that do not satisfy $\cal C$. The condition $V_{\phi}(\hat{Y})\leq 0\cup\Delta V_{\phi}(\hat{Y})\geq 0$ restricts the found state $\hat{y}$ from satisfying the conditions in Equation 7. The constraint $\lVert\hat{Y_{i}}-Y_{i}\rVert_{2}\geq\epsilon$ ensures numerical stability by excluding states too close to equilibrium, which may otherwise introduce numerical instability during the SMT solving process. It is important to emphasize that these counterexamples are not limited to predictions generated from the original well-trained GNN. Instead, we employ an SMT solver to systematically explore a broader space of node prediction states, intentionally enrich the training set for training the neural controller $f_{\theta}$ and the neural Lyapunov function $V_{\phi}$. Our training process is detailed in Algorithm 1.
@@
16:     Prepare closed-loop system $C$ and $\hat{Y}$ derived from well-trained GNN
@@
III-D Test-Time Feature Reconstruction
@@
It is important to note that we do not need to perform multiple iterations during testing. The reason is that the trained neural controller and neural Lyapunov function have already stabilized the control system. When the system is at the equilibrium point, feeding this state back into the controller results in control inputs that do not alter the equilibrium point. Therefore, a single adjustment suffices to bring the node features to stable class representative embeddings, and further iterations are unnecessary.
TABLE II: Results(%) for the semi-supervised node classification task on the four datasets. Training split represents the method for selecting training nodes. ’Bias’ means that training nodes are selected using the Scalable Biased Sampler [19]. Biased training samples simulate the situation with distribution shift. Each result is reported as an average $\pm$ standard deviation across 10 experiments. The footnote of FRGNN indicates the base model of FRGNN. * indicates that the FRGNN only replaces one class of node features. OOM indicates that the method runs out of memory. The best results are highlighted in bold and the second-best results are underlined.
@@ IV Experiments
IV-A Experimental Setup
IV-A1 Base Model
We chose SGC  [20] as our base model. SGC is a simple graph neural network that reduces the multi-layer convolution operations of GCN  [21] to single matrix multiplication, making it more efficient. Our method requires using an SMT solver to identify states that violate Lyapunov constraints, so we must keep the SMT solving time within an acceptable range. Thus, we select SGC as the base model to accelerate the solving process by reducing the graph neural network’s parameters.
IV-A2 Datasets
We evaluate our method on five popular graph datasets: Cora, Citeseer, Pubmed [22], Amazon Photo, and Amazon Computers [23]. These datasets are standard benchmarks for semi-supervised node classification tasks. The statistics of these datasets are summarized in Appendix B. For each dataset, we follow the setup from FRGNN  [10] and apply the Scalable Biased Sampler [19] to create training, validation, and test sets. This sampling method effectively captures datasets with distribution shifts, allowing us to demonstrate the ability of our method to address such issues.
TABLE III: Classification accuracy comparison between our neural controller and traditional controller on three datasets. Each result is reported as an average $\pm$ standard deviation across 10 experiments. MFAC indicates that the controller is MFAC and NC indicates that the controller is our neural controller. The best results are highlighted in bold.
@@ IV-A2 Datasets
IV-A3 Implementation Details
Our training process for the neural controller and the neural Lyapunov function heavily relies on generating counterexamples, which necessitates ensuring the speed of the SMT solver. Since the speed of SMT solver is proportional to the number of parameters in each graph neural network layer, we downsample the node features to reduce the parameter count. We apply PCA to reduce the feature dimensions to a fixed 20 dimensions, making it the input size for all experiments. We use SGC as the base model, setting its steps of feature propagation to 3 and its output size to the number of classes. Our neural controller $f_{\theta}$ is a two-layer MLP with a hidden size of 16. Our neural Lyapunov function $V_{\phi}$ is also a two-layer MLP with a hidden size of 16 and an output size of 1. We train both $f_{\theta}$ and $V_{\phi}$ using the Adam optimizer with a learning rate of 0.001. We conduct all experiments on an RTX 3090 24G GPU.
During the training phase, we randomly select a node as the initial node and use the one-hot embedding of its class as the equilibrium point $Y$. We use the initial prediction $\hat{Y}$ of this node as the initial state. We then adjust the node features through the neural controller $f_{\theta}$ and input the reconstructed node features into the well-trained GNN to obtain the next prediction $\hat{Y}_{t+1}$. With this information, we calculate the Lyapunov loss and update the neural controller $f_{\theta}$ and the neural Lyapunov function $V_{\phi}$. Then we use the SMT solver to find counterexamples that violate the Lyapunov criterion. When the solver finds such states within the domain, we add these states as counterexamples to the training set. We repeat this process until all states within the domain satisfy the Lyapunov stability criterion. At this point, we obtain a neural controller $f_{\theta}$ that satisfies the Lyapunov stability criterion.
During the testing phase, we input the one-hot embedding of the selected node into the neural controller $f_{\theta}$ to obtain the reconstructed node features. These reconstructed features are then used as class representative embeddings for the selected class, replacing the features of labeled nodes within that class. Finally, we input these updated node features into the well-trained GNN to assess the model’s performance on the test set.
Remark.
In this work, we focus specifically on reconstructing node features belonging to a single node class, treating the ground truth labels of this class as the sole equilibrium point in our Lyapunov stability framework. While extending this method to multiple classes (with multiple equilibrium points) is theoretically straightforward, such an extension significantly increases the complexity of the SMT solving process in Algorithm  Algorithm 1. Investigating efficient SMT solver to handle multiple equilibrium points requires future research.
@@ IV-A3 Implementation Details
IV-A4 Baseline
We select MH-AUG  [24], KDGA  [25], and FRGNN  [10] as the baselines. More detail about these methods is provided in the Appendix C.
@@ IV-A4 Baseline
Remark.
Our chosen baselines include both methods that directly reconstruct node features without modifying model parameters or structure (e.g., FRGNN), as well as more sophisticated methods that utilize data augmentation, knowledge distillation, or other techniques (e.g., MH-AUG, KDGA). The motivation for selecting such diverse baselines is to comprehensively illustrate the fundamental differences between our proposed control-theoretic approach and existing mainstream methods designed to mitigate distribution shifts from various perspectives.
However, due to computational constraints introduced by the SMT solver, our experimental validation is necessarily conducted on relatively small datasets. Like many other works  [26, 27, 28], our primary objective is not to outperform existing state-of-the-art graph neural networks on standard datasets, but rather to validate the theoretical correctness and practical effectiveness of our novel approach of applying our novel framework to existing GNNs, specifically the application of control theory and neural Lyapunov stability analysis to GNN feature reconstruction problems under distribution shifts.
IV-B Results
IV-B1 Performance Comparison
@@ V Conclusion
Appendix A
Theorem 1: Consider a closed-loop system $\hat{Y}_{t+1}=C_{f_{\theta}}(\hat{Y_{t}})$ with an equilibrium point at the $Y$, i.e. $C_{f_{\theta}}(Y)=Y$, where $C_{f_{\theta}}$ is a MLP with ReLU activation function. Suppose there exists a continuously differentiable function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ that satisfies the following conditions:
$\Delta V=V(C_{f_{\theta}}(\hat{Y_{t}}))-V(\hat{Y_{t}}).$ (13)
Then, the system is asymptotically stable at the $Y$ and $V$ is called a Lyapunov function. theorem
Proof.
We consider the neural network controller $f_{\theta}$ and the classifier $C$ as an MLP $C_{f_{\theta}}$ with $l+k$ layers. For an MLP with $l+k$ layers, we have:
$C^{l+k}_{f_{\theta}}(x)$ $=g^{(l+k)}(g^{(l+k-1)}(\cdots g^{(1)}(x)\cdots)),$
where $g^{(i)}$ represents the $i$-th layer of the MLP. For a ReLU activation function, $g^{i}$ can be expressed as:
$g^{(i)}(x)=ReLU(W^{(i)}x+b^{(i)}),$
For any $x_{1},x_{2}\in\mathbb{R}^{n}$, we have:
$\lVert C^{l+k}_{f_{\theta}}(x_{1})-C^{l+k}_{f_{\theta}}(x_{2})\rVert_{2}$
$=\lVert g^{(l+k)}(g^{(l+k-1)}(\cdots g^{(1)}(x_{1})\cdots))-$
$g^{(l+k)}(g^{(l+k-1)}(\cdots g^{(1)}(x_{2})\cdots))\rVert_{2}$
$\leq\lVert W^{l+k}g^{(l+k-1)}(\cdots g^{(1)}(x_{1})\cdots)+B^{l+k}$
$-W^{l+k}g^{(l+k-1)}(\cdots g^{(1)}(x_{2})\cdots)-B^{l+k}\rVert_{2}$
[…]
$\leq\lVert g^{(l+k-1)}(\cdots g^{(1)}(x_{1})\cdots)-$
$g^{(l+k-1)}(\cdots g^{(1)}(x_{2})\cdots)\rVert_{2}\cdot\lVert W^{(l+k)}\rVert_{2}$
[…]
$\leq\lVert x_{1}-x_{2}\rVert_{2}\cdot\prod_{i=1}^{l+k}\lVert W^{(i)}\rVert_{2}$
$=L\lVert x_{1}-x_{2}\rVert_{2}.$
Therefore, $C_{f_{\theta}}$ is locally Lipschitz continuous. For a locally Lipschitz continuous function, we can apply the Lyapunov stability criterion. Refer to  [12], we rewritten the Lyapunov stability criterion as follows:
$V(x_{0})=0,and,\forall x\in\mathcal{D}\setminus\{x_{0}\},V(x)>0\>and\>\Delta V\leq 0,$
where $x_{0}$ is the equilibrium point, $V$ is the Lyapunov function, $D$ is the domain of the control system, and $\Delta V$ is defined as:
$\Delta V=V(F(x))-V(x),$
where $F$ is locally Lipschitz continuous.We replace $F$ with $C_{f_{\theta}}$ and $x$ with $\hat{Y_{t}}$, then we have:
$V(\hat{Y})=0,and,\forall\hat{Y}\in\mathcal{D}\setminus\{Y\},V(\hat{Y})>0\>and\>\Delta V\leq 0.$
When the above conditions are satisfied, the system is asymptotically stable at the equilibrium point $Y$. This completes the proof.
TABLE IV: Overall datasets statistics
Dataset Cora Citeseer Pubmed Amazon Photo Amazon Computers
#Nodes 2,708 3,327 19,717 7487 13381
#Edges 5,429 4,732 44,338 119043 245778
#Features 1,433 3,703 500 745 767
#Classes 7 6 3 8 10
#Label Rate 5.2% 3.6% 0.3% 2.14% 1.49%
#Train Nodes 140 120 60 160 200
#Val Nodes 500 500 500 240 300
#Test Nodes 1,000 1,000 1,000 7250 13252
Appendix B
We summarize the statistics of the datasets used in our experiments in LABEL:tab2. The datasets include Cora, Citeseer, Pubmed, Amazon Photo, and Amazon Computers. Each dataset contains a different number of nodes, edges, features, and classes. The label rate of each dataset is also provided.
Appendix C
MH-AUG improves model generalization through adaptive data augmentation. KDGA enhances the model’s adaptability to distribution shift by training a teacher model with augmented data and distilling its knowledge into a student model. FRGNN utilizes the predictions of a well-trained GNN and node features to train an MLP, which generates class representative embeddings using the one-hot embedding of each class. These features are then used to replace the features of labeled nodes, improving robustness against distribution shift.